OpenStudy (anonymous):

c Cathy spent 3/4 of her pocket money. If she spent 1/3 of it for a movie, 1/6 for books and the rest for a dress, what part of her pocket money ...

OpenStudy (anonymous):

Let \(c\) be the amount of pocket money Cathy has, and \(x\) be the fraction she spent on the dress. Then\[\frac{3}{4}c=\frac{1}{3}c+\frac{1}{6}c+xc\]Can you take it from there?

OpenStudy (anonymous):

pls solve

OpenStudy (anonymous):

No, YOU solve.

OpenStudy (anonymous):

pls help me to solve

OpenStudy (anonymous):

pls show me the first step

OpenStudy (anonymous):

Cancel out all of the \(c\) terms.

OpenStudy (anonymous):

*\(c\) factors

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

i did nt get

OpenStudy (anonymous):

my answer is 1/4, but in my maths text book answer is 3/8